# Stop Overthinking: A Survey on Efficient Reasoning for Large Language Models

### v2

@@ 1 Introduction
In this paper, we present the first structured survey systematically exploring the progress in efficient reasoning for LLMs. As illustrated in Figure 2, we categorize existing work into three key directions: (1) Model-based efficient reasoning, which focuses on optimizing full-length reasoning models into more concise variants or directly training efficient reasoning models. (2) Reasoning output-based efficient reasoning, which dynamically reduces reasoning steps and length during inference. (3) Input prompts-based efficient reasoning, which enhances reasoning efficiency based on input properties such as difficulty or length control. Unlike model compression techniques such as quantization [97, 29, 50] or KV cache compression [110, 55, 76, 107], which focus on reducing model size for lightweight inference, efficient reasoning in LLMs emphasizes smart and concise reasoning by optimizing the length of generated reasoning sequences and reducing unnecessary thinking steps.
Overall, we provide a summary of the current key approaches to efficient reasoning, organizing them into following categories:
@@
, ERTleaf node] ] ] [ Reasoning Output-based Efficient Reasoning , OER node [ Latent Representation Compression , OER node [e.g. Coconut [35]; CODI [75]; CCoT [12]; Heima [73]; Token Assorted [79]; Loop [70]; SoftCoT [100];
, OERleaf node] ] [ Dynamic Reasoning Paradigm , OER node [e.g. Speculative Rejection [81]; Sampling-Efficient TTS [91]; DPTS [26]; Certaindex [30]; Fast MCTS [44]; ST-BoN [91]; More is Less [94]; RSD [48]; LightThinker [108]; INFTYTHINK [101]; SCoT [96]; RASC [89]; Adaptive Reasoning [106]; AdaptiveStep [54]; , OERleaf node] ] ] [ Input Prompts-based Efficient Reasoning , ERI node [ Prompt-Guided Efficient Reasoning , ERI node [e.g. Token-Budget [34]; Chain of Draft [99]; Token Complexity [43]; Concise Chain-of-Thought (CCoT) [69]; MARP [8]; , ERIleaf node] ] [ Routing by Question Attributes , ERI node [e.g. Claude 3.7 Sonnet [2]; SoT [4]; Self-REF [17]; Confident [14]; RouteLLM [64]; , ERIleaf node] ] ] [ Efficient Data and Models , CP node [ Less Training Data , CP node [e.g. LIMO [103]; s1 [61]; S2[58]; Light-R1 [93]; , CPleaf node] ] [ Pruning & Quantization & Distillation , CP node [e.g. Struggle[46]; Strong Verifiers [78]; TinyR1-32B-Preview [82]; Mixed Distillation [13]; Counterfactual Distillation [28]; Feedback-Driven Distillation [112]; SKIntern [49]; Adaptive Thinking [9]; PRR [111]; , CPleaf node] ] ] [ Benchmark & Insights , BI node [ Evaluation & Benchmarks , BI node [e.g. 1B vs. 405B [52]; Sys2Bench [66]; Danger [21]; Inference-time Computation [51]; Impact [39]; , BIleaf node] ] ] ]
@@ 3.1 RL with Length Reward Design
Table 2: Comparison of different policy optimization methods in CoT length controls. $\hat{R}_{t}$ represents the reward model. $\pi_{\text{ref}}$ is the policy of reference model. $\gamma$ is a target reward margin term for SimPO. The $y_{w}$ is for winning responses and $y_{l}$ is for losing responses, where some with $G$ on superscript denote the outputs of different sampled groups.
@@
PPO [71] $\mathbb{E}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y_{w}\mid x)}{\pi_{\text{ref}}(y_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y_{l}\mid x)}{\pi_{\text{ref}}(y_{l}\mid x)}\right)\right]$
SimPO [60] $\mathbb{E}\left[\log\sigma\left(\frac{\beta}{|y_{w}|}\log\pi_{\theta}(y_{w}\mid x)-\frac{\beta}{|y_{l}|}\log\pi_{\theta}(y_{l}\mid x)-\gamma\right)\right]$
GRPO [72] $\mathbb{E}\left[\log\sigma\left(\beta\log\frac{\pi_{\theta}(y^{G}_{w}\mid x)}{\pi_{\text{ref}}(y^{G}_{w}\mid x)}-\beta\log\frac{\pi_{\theta}(y^{G}_{l}\mid x)}{\pi_{\text{ref}}(y^{G}_{l}\mid x)}\right)-\lambda\mathbb{D}_{KL}[\pi_{\theta}||\pi_{\text{ref}}]\right]$
@@ 3.2 SFT with Variable-Length CoT Data
CoT-Value [59] GSM8K PRM800k Merging parameters of non-reasoning and long reasoning LLMs Progressive QwQ-32B-Preview DeepSeek-R1-Distill-Llama-8B LLaMA-3.1-8B LLaMA-3.2-1B Qwen32B-Instruct
@@ 3.2.1 Constructing Variable-Length CoT Reasoning Datasets
This approach collects short CoT data by prompting LLMs to generate short reasoning steps during inference and reasoning, as proposed in [53], [62][34], and [59]. Specifically, [53] proposes a human-like step-skipping method for generating shorter reasoning steps. In the first stage, based on the original training datasets, they manually create solutions by skipping steps, either guided by human expertise or by randomly merging or removing steps. Further, these concise data are labeled with prompts such as “Solve it in $n$ steps.”. After SFT, the model is able to generate shorter reasoning paths. In the second stage, they prompt this model to solve problems by intrinsically skipping or compressing steps during reasoning. The generated concise reasoning steps with questions and answers are collected as datasets, which are then used in SFT to make LLMs solve problems with fewer steps. Moreover, Token-Budget [34] has an important insight: an optimal token budget helps LLMs actively follow the token constraint to complete the reasoning process. Motivated by this insight, it proposes a binary search-based method to achieve the optimal token budgets, and follow these budgets to generate short reasoning steps. In addition, [62] proposes a sampling-based method to improve reasoning efficiency. Specifically, it examines the distribution of reasoning lengths and finds that shorter solutions appear more frequently than the typical reasoning length. Driven by this finding, it proposes a Best-of-N (BoN) Sampling at test time, which generates $N$ paths of reasoning and selects the shortest one. These short reasoning paths are collected as the dataset. Finally, CoT-Value [59] controls the reasoning length by mix-up the parameters of long reasoning and non-reasoning LLMs for generating variable-length reasoning steps. The advantage of CoT compression during reasoning is that the naturally generated reasoning steps align with the intrinsic knowledge of LLMs, which advances more effective learning of LLMs.
@@ 3.2.2 Fine-Tuning Approaches
Progressive fine-tuning aims to smoothly reduce the reasoning steps during fine-tuning [59, 53]. One way is to progressively reduce the reasoning steps of data during fine-tuning LLMs, as employed in [53]. Another effective way is to progressively adjust the generation of reasoning steps, as proposed by CoT-Value [59]. Specifically, it first learns LoRA adaptor $\Delta\theta_{\text{N}}$ and $\theta_{\text{L}}$, where LLMs with $\Delta\theta_{\text{N}}$ have no reasoning steps, and that with $\Delta\theta_{\text{L}}$ have long reasoning. Then, it mix-up $\Delta\theta_{\text{N}}$ and $\Delta\theta_{\text{long}}$ by $\alpha\Delta\theta_{\text{N}}+(1-\alpha)\Delta\theta_{\text{L}}$ to generate a dataset reasoning with variable length. Here $10<\alpha<$ 1 controls the parameter to shift from $\Delta\theta_{\text{N}}$ to $\Delta\theta_{\text{L}}$, controlling the reasoning length generated by LLMs. Finally, it fine-tunes LLMs on the generated data while progressively reducing $\alpha$ from 1 to 0. In this way, it progressively improves the reasoning efficiency during fine-tuning.
@@ 4.2.1 Dynamic Reasoning via Explicit Criteria
Dynamic Parallel Tree Search (DPTS) [26] optimizes tree-based reasoning in LLMs by addressing two main inefficiencies by introducing: (1) Parallelism Streamline optimizes memory and compute by storing only incremental KV cache updates and dynamically adjusting the number of extended nodes based on available GPU memory, (2) Search and Transition Mechanism balances exploration and exploitation using confidence-based criteria. Overall, during inference, the system cuts off uncertain paths to save time. FastMCTS [44] is another confidence-based method that aims to optimize multi-step reasoning data synthesis. Traditional rejection sampling generates multiple candidate responses independently, selecting only the correct ones, but it is often inefficient and struggles with imbalanced sampling. Inspired by MCTS, FastMCTS prioritizes high-confidence traces for deep reasoning. Additionally, it adjusts tree expansion based on problem complexity, improving both efficiency and reasoning diversity. Another line of research leverages certainty or uncertainty measures to guide adaptive reasoning. Certaindex [30], a certainty metric, quantifies the confidence of LLMs throughout reasoning using semantic entropy, reward model scores, or a combination of both. A higher certaindex indicates that further reasoning steps are unlikely to change the final answer, allowing early termination to free resources for more challenging queries. Dynasor, an inference system built on this principle, optimizes compute scheduling by dynamically tracking reasoning progress instead of allocating resources uniformly. Length-filtered Vote [94] is another work that leverages uncertainty to improve CoT reasoning. The study finds that longer reasoning chains do not always improve accuracy; instead, performance initially improves but eventually declines due to error accumulation. The authors provide a mathematical analysis proving the existence of an optimal CoT length, determined by model capability and task difficulty. To exploit this, they propose Length-filtered Voting, a length-aware majority voting method that groups answers by CoT length and selects the most reliable group based on prediction uncertainty.
@@ 7 Evaluation and Benchmark
[66] develops Sys2Bench, which is a comprehensive suite designed to evaluate LLMs across five reasoning categories, including arithmetic, logical, commonsense, algorithmic, and planning tasks. This benchmark comprises eleven diverse datasets, covering various reasoning tasks. It includes GSM8K and AQuA for arithmetic problems, StrategyQA and HotPotQA for commonsense reasoning, ProntoQA for logical reasoning, Game of 24 and Bin Packing for algorithmic tasks, and BlocksWorld, Rubik’s Cube, TripPlan, and Calendar Plan for planning tasks. The study revealed that scaling inference-time computation alone has limitations, as no single technique consistently excels across all reasoning tasks, and this emphasizes the need for diverse approaches to enhance LLM reasoning capabilities. [51] examine how various commonly used strategies affect the reasoning capabilities of LLMs. Further, they present an extensive experimental benchmark involving six inference-time optimization techniques across eight reasoning-oriented tasks. [52] investigates the impact of Test-Time Scaling (TTS) strategies on LLM performance, focusing on how policy models, process reward models, and problem difficulty influence TTS effectiveness. Their findings indicate that compute-optimal TTS strategies are highly dependent on these factors. The paper finds that, with appropriate TTS strategies, smaller models (e.g., a 1B parameter LLM) are able to outperform significantly larger models (e.g., a 405B parameter LLM) on complex reasoning tasks like MATH-500, and this underscores the importance of tailored TTS approaches in evaluating and enhancing LLM reasoning.
@@
[21] introduces a framework to systematically analyze the "overthinking" in LLMs, where models favor extended internal reasoning over necessary environmental interactions. By examining 4,018 trajectories in agentic tasks, the study identified patterns such as Analysis Paralysis, Rogue Actions, and Premature Disengagement. [21] also proposed a novel “overthinking score” and showed a strong correlation between higher scores and decreased task performance. Mitigation strategies such as selecting solutions with lower overthinking scores can improve performance by 30% and at the same time reduce computational overhead by 43%.
@@
[104] provides a comprehensive analysis of the mechanism underlying long CoT reasoning. In addition to presenting several key insights, they propose a reward design to enhance the stability of reasoning ability during training and reduce the CoT length, which is also shown in Section 3.1. [39] reveals a strong relationship between the length of the reasoning chain and the effectiveness of model outputs. Models tend to perform better with extended reasoning steps, suggesting the CoT length is more crucial than accuracy for effective problem-solving.
@@ 8.1 Applications
Efficient reasoning LLMs are able to greatly improve autonomic driving [22] by helping them understand large amounts of sensor data in a human-like way. They make the cars better at making decisions, so the vehicles can plan for difficult driving situations and react quickly when unexpected events occur. By combining information from cameras, LiDAR, radar, and other sensors, these models help cars drive more safely, choose better routes, and assess risks as they happen. Moreover, because they can explain why they make certain decisions, both passengers and regulators feel more confident in the technology, and the cars can interact more smoothly with smart road systems.

### v3

@@ 1 Introduction
In this paper, we present the first structured survey systematically exploring the progress in efficient reasoning for LLMs. As illustrated in Figure 2, we categorize existing work into three key directions: (1) Model-based efficient reasoning, which focuses on optimizing full-length reasoning models into more concise variants or directly training efficient reasoning models. (2) Reasoning output-based efficient reasoning, which dynamically reduces reasoning steps and length during inference. (3) Input prompts-based efficient reasoning, which enhances reasoning efficiency based on input properties such as difficulty or length control. Unlike model compression techniques such as quantization [126, 31, 62] or KV cache compression [150, 73, 100, 145, 38], which focus on reducing model size for lightweight inference, efficient reasoning in LLMs emphasizes smart and concise reasoning by optimizing the length of generated reasoning sequences and reducing unnecessary thinking steps.
Overall, we provide a summary of the current key approaches to efficient reasoning, organizing them into the following categories:
@@ 3.1 RL with Length Reward Design
Table 2: Comparison of different policy optimization methods in CoT length controls. $\hat{R}_{t}$ represents the reward model. $\pi_{\text{ref}}$ is the policy of reference model. $\gamma$ is a target reward margin term for SimPO. $\lambda$ is a clipping-related hyper-parameter. The $y_{w}$ is for winning responses, and $y_{l}$ is for losing responses, where some with $G$ on superscript denote the outputs of different sampled groups.
@@
PPO [94] $\mathbb{E}\left[\min\Big{(}\frac{\pi_{\theta}(y_{t}\mid x_{t})}{\pi_{\theta_{\text{ref}}}(y_{t}\mid x_{t})}\hat{R}_{t},\ \text{clip}\left(\frac{\pi_{\theta}(y_{t}\mid x_{t})}{\pi_{\theta_{\text{ref}}}(y_{t}\mid x_{t})},\ 1-\epsilon,\ 1+\epsilon\right)\hat{R}_{t}\Big{)}\right]$
SimPO [80] $\mathbb{E}\left[\log\sigma\left(\frac{\beta}{|y_{t}^{w}|}\log\pi_{\theta}(y_{t}^{w}\mid x_{t})-\frac{\beta}{|y_{t}^{l}|}\log\pi_{\theta}(y_{t}^{l}\mid x_{t})-\gamma\right)\right]$
GRPO [95] $\mathbb{E}\left[\min\Big{(}\frac{\pi_{\theta}(y^{G}_{t}\mid x_{t})}{\pi_{\theta_{\text{ref}}}(y^{G}_{t}\mid x_{t})}\hat{R}^{G}_{t},\ \text{clip}\left(\frac{\pi_{\theta}(y^{G}_{t}\mid x_{t})}{\pi_{\theta_{\text{r}}}(y^{G}_{t}\mid x_{t})},\ 1-\epsilon,\ 1+\epsilon\right)\hat{R}^{G}_{t}\Big{)}-\lambda\mathbb{D}_{KL}[\pi_{\theta}||\pi_{\text{ref}}]\right]$
@@ 3.2 SFT with Variable-Length CoT Data
CoT-Valve [79] GSM8K PRM800k Merging parameters of non-reasoning and long reasoning LLMs Progressive QwQ-32B-Preview DeepSeek-R1-Distill-Llama-8B LLaMA-3.1-8B LLaMA-3.2-1B Qwen32B-Instruct
@@ 3.2.1 Constructing Variable-Length CoT Reasoning Datasets
This approach collects short CoT data by prompting LLMs to generate short reasoning steps during inference and reasoning, as proposed in [70], [82][37], and [79]. Specifically, [70] proposes a human-like step-skipping method for generating shorter reasoning steps. In the first stage, based on the original training datasets, they manually create solutions by skipping steps, either guided by human expertise or by randomly merging or removing steps. Further, these concise data are labeled with prompts such as “Solve it in $n$ steps.”. After SFT, the model is able to generate shorter reasoning paths. In the second stage, they prompt this model to solve problems by intrinsically skipping or compressing steps during reasoning. The generated concise reasoning steps with questions and answers are collected as datasets, which are then used in SFT to make LLMs solve problems with fewer steps. Moreover, Token-Budget [37] has an important insight: an optimal token budget helps LLMs actively follow the token constraint to complete the reasoning process. Motivated by this insight, it proposes a binary search-based method to achieve the optimal token budgets, and follow these budgets to generate short reasoning steps. In addition, [82] proposes a sampling-based method to improve reasoning efficiency. Specifically, it examines the distribution of reasoning lengths and finds that shorter solutions appear more frequently than the typical reasoning length. Driven by this finding, it proposes a Best-of-N (BoN) Sampling at test time, which generates $N$ paths of reasoning and selects the shortest one. These short reasoning paths are collected as the dataset. Finally, CoT-Valve [79] controls the reasoning length by mix-up the parameters of long reasoning and non-reasoning LLMs for generating variable-length reasoning steps. The advantage of CoT compression during reasoning is that the naturally generated reasoning steps align with the intrinsic knowledge of LLMs, which advances more effective learning of LLMs.
@@ 3.2.2 Fine-Tuning Approaches
Progressive fine-tuning aims to smoothly reduce the reasoning steps during fine-tuning [79, 70]. One way is to progressively reduce the reasoning steps of data during fine-tuning LLMs, as employed in [70]. Another effective way is to progressively adjust the generation of reasoning steps, as proposed by CoT-Valve [79]. Specifically, it first learns LoRA adaptor $\Delta\theta_{\text{N}}$ and $\theta_{\text{L}}$, where LLMs with $\Delta\theta_{\text{N}}$ have no reasoning steps, and that with $\Delta\theta_{\text{L}}$ have long reasoning. Then, it mix-up $\Delta\theta_{\text{N}}$ and $\Delta\theta_{\text{long}}$ by $\alpha\Delta\theta_{\text{N}}+(1-\alpha)\Delta\theta_{\text{L}}$ to generate a dataset reasoning with variable length. Here $10<\alpha<$ 1 controls the parameter to shift from $\Delta\theta_{\text{N}}$ to $\Delta\theta_{\text{L}}$, controlling the reasoning length generated by LLMs. Finally, it fine-tunes LLMs on the generated data while progressively reducing $\alpha$ from 1 to 0. In this way, reasoning efficiency is progressively improved during fine-tuning.
Model Merging.
Beyond fine-tuning reasoning models using the above CoT data, several works have explored model merging to improve reasoning efficiency in LLMs. Kimi k1.5 [109] investigates merging strategies [135] that transform models producing lengthy CoT traces into ones that generate shorter, more concise reasoning outputs. Unlocking [122] conducts an empirical study on model merging for efficient reasoning with various methods, including task-vector-based, SVD-based, and activation-informed merging.
@@ 4.2.1 Dynamic Reasoning via Explicit Criteria
Dynamic Parallel Tree Search (DPTS) [25] optimizes tree-based reasoning in LLMs by addressing two main inefficiencies by introducing: (1) Parallelism Streamline optimizes memory and compute by storing only incremental KV cache updates and dynamically adjusting the number of extended nodes based on available GPU memory, (2) Search and Transition Mechanism balances exploration and exploitation using confidence-based criteria. Overall, during inference, the system cuts off uncertain paths to save time. FastMCTS [55] is another confidence-based method that aims to optimize multi-step reasoning data synthesis. Traditional rejection sampling generates multiple candidate responses independently, selecting only the correct ones, but it is often inefficient and struggles with imbalanced sampling. Inspired by MCTS, FastMCTS prioritizes high-confidence traces for deep reasoning. Additionally, it adjusts tree expansion based on problem complexity, improving both efficiency and reasoning diversity. Another line of research leverages certainty or uncertainty measures to guide adaptive reasoning. Certaindex [32], a certainty metric, quantifies the confidence of LLMs throughout reasoning using semantic entropy, reward model scores, or a combination of both. A higher certaindex indicates that further reasoning steps are unlikely to change the final answer, allowing early termination to free resources for more challenging queries. Dynasor, an inference system built on this principle, optimizes compute scheduling by dynamically tracking reasoning progress instead of allocating resources uniformly. Dynasor-CoT [33] uses a probing scheme to exit early if the model is confident enough. Length-filtered Vote [123] is another work that leverages uncertainty to improve CoT reasoning. The study finds that longer reasoning chains do not always improve accuracy; instead, performance initially improves but eventually declines due to error accumulation. The authors provide a mathematical analysis proving the existence of an optimal CoT length, determined by model capability and task difficulty. To exploit this, they propose Length-filtered Voting, a length-aware majority voting method that groups answers by CoT length and selects the most reliable group based on prediction uncertainty. Self-Calib [46] introduces confidence-driven adaptive scaling strategies at test time to efficiently address queries with differing complexity levels, including Early-Stopping mechanisms for Best-of-N sampling and confidence-calibrated Self-Consistency approaches. CISC [108] implements a weighted majority voting scheme utilizing model-derived confidence scores. By emphasizing reasoning paths with higher confidence, it efficiently determines the correct response with substantially fewer samples.
@@ 7 Evaluation and Benchmark
[88] develops Sys2Bench, which is a comprehensive suite designed to evaluate LLMs across five reasoning categories, including arithmetic, logical, commonsense, algorithmic, and planning tasks. This benchmark comprises eleven diverse datasets, covering various reasoning tasks. It includes GSM8K and AQuA for arithmetic problems, StrategyQA and HotPotQA for commonsense reasoning, ProntoQA for logical reasoning, Game of 24 and Bin Packing for algorithmic tasks, and BlocksWorld, Rubik’s Cube, TripPlan, and Calendar Plan for planning tasks. The study revealed that scaling inference-time computation alone has limitations, as no single technique consistently excels across all reasoning tasks, and this emphasizes the need for diverse approaches to enhance LLM reasoning capabilities. [65] examine how various commonly used strategies affect the reasoning capabilities of LLMs. Further, they present an extensive experimental benchmark involving six inference-time optimization techniques across eight reasoning-oriented tasks. [68] investigates the impact of Test-Time Scaling (TTS) strategies on LLM performance, focusing on how policy models, process reward models, and problem difficulty influence TTS effectiveness. Their findings indicate that compute-optimal TTS strategies are highly dependent on these factors. The paper finds that, with appropriate TTS strategies, smaller models (e.g., a 1B parameter LLM) are able to outperform significantly larger models (e.g., a 405B parameter LLM) on complex reasoning tasks like MATH-500, and this underscores the importance of tailored TTS approaches in evaluating and enhancing LLM reasoning. Bag of Tricks [66] investigates several often-overlooked techniques capable of enhancing the reasoning performance of LLMs. Furthermore, it benchmarks multiple inference-time computation methods within a predefined budget, enabling controlled token usage through a flexible N-sample strategy.
@@
[20] introduces a framework to systematically analyze the "overthinking" in LLMs, where models favor extended internal reasoning over necessary environmental interactions. By examining 4,018 trajectories in agentic tasks, the study identified patterns such as Analysis Paralysis, Rogue Actions, and Premature Disengagement. [20] also proposed a novel “overthinking score” and showed a strong correlation between higher scores and decreased task performance. Mitigation strategies such as selecting solutions with lower overthinking scores can improve performance by 30% and at the same time reduce computational overhead by 43%. [69] uses six comprehensive tasks from the overthinking literature to guide the design of evaluation benchmarks for testing CoT failures in LLMs. It reveals that, in many cases where humans tend to fail due to excessive deliberation, LLMs employing CoT reasoning exhibit similar failure patterns. S1-Bench [148] evaluates large reasoning models on straightforward tasks aligned with System 1 thinking, focusing on intuitive and fast reasoning rather than the more complex, deliberative processes associated with System 2.
@@
[141] provides a comprehensive analysis of the mechanism underlying long CoT reasoning. In addition to presenting several key insights, they propose a reward design to enhance the stability of reasoning ability during training and reduce the CoT length, which is also shown in Section 3.1. [47] reveals a strong relationship between the length of the reasoning chain and the effectiveness of model outputs. Models tend to perform better with extended reasoning steps, suggesting the CoT length is more crucial than accuracy for effective problem-solving. CriticalThinking [53] investigates the optimal reasoning length of LLMs based on using deterministic finite automata (DFAs). MiP-Overthinking [28] discovers that when the reasoning models get questions missing key information, they tend to keep “thinking” over and over instead of admitting they fail to answer. By studying this behavior, they show that current training methods push models to generate long, repetitive reasoning steps rather than recognize unsolvable problems and stop early.
Effect of Compression on Reasoning Models.
CompressionReasoning [147] benchmarks compression techniques including quantization, distillation, and pruning, on reasoning tasks. The results indicate that parameter count has a greater impact on knowledge retention than on reasoning ability, and that generating shorter outputs generally leads to improved performance. QuantRM [67] provides a benchmark evaluating weight quantization, KV-cache quantization, and activation quantization across various algorithms and bit-width configurations.
@@ 8.1 Applications
Efficient reasoning LLMs are able to greatly improve autonomic driving [21, 128, 127, 129] by helping them understand large amounts of sensor data in a human-like way. They make the cars better at making decisions, so the vehicles can plan for difficult driving situations and react quickly when unexpected events occur. By combining information from cameras, LiDAR, radar, and other sensors, these models help cars drive more safely, choose better routes, and assess risks as they happen. Moreover, because they can explain why they make certain decisions, both passengers and regulators feel more confident in the technology, and the cars can interact more smoothly with smart road systems.
@@
Recommender System.
Efficient reasoning LLMs can greatly enhance recommender systems by enabling more accurate, personalized, and context-aware suggestions across various domains such as e-commerce, entertainment, and education. These models can reason over diverse and dynamic user behavior, preferences, and historical interactions to uncover subtle patterns and relationships that traditional models might overlook. By efficiently processing complex input data, LLMs can generate high-quality recommendations with fewer computational resources. For instance, in an online shopping platform, an efficient reasoning model can anticipate evolving user interests, adapt to seasonal trends, and explain recommendations in a user-friendly way. Overall, efficient reasoning LLMs improve the scalability, transparency, and responsiveness of recommender systems, leading to better user satisfaction and engagement. ReaRec [107] proposes a latent reasoning framework for recommender systems. Inspired by the think-before-action paradigm in LLMs, ReaRec enables implicit multi-step reasoning at inference time, significantly improving performance, particularly for long-tail users and items.